# Uni-3DAR: Unified 3D Generation and Understanding via Autoregression on Compressed Spatial Tokens

### v2

@@ 2.3 Unified 3D Generation and Understanding Framework
With the two optimizations described above, Uni-3DAR reduces the token count by approximately 8x while maintaining strong performance. Additionally, by leveraging duplicated masked tokens, Uni-3DAR can unify a broad range of existing 3D structural tasks within a single model.
Specifically, 3D structural generation and understanding tasks can be categorized into four types, which may be applied individually or in combination:
@@
    Single-Frame Generation: This task processes a single 3D structure. Generation may be unconditional or conditioned on external modalities such as properties or textual descriptions. In Uni-3DAR, generation is performed on masked tokens, making this the most straightforward application.
@@
    Multi-Frame Generation: This task involves generating a sequence of 3D structures, typically conditioned on an initial structure. Common applications include molecular dynamics, where an initial conformation conditions the generation of subsequent states, and pocket-based generation, where a 3D protein pocket guides the synthesis of a binding molecule. Leveraging Uni-3DAR’s autoregressive generation, structures are produced frame by frame, with each frame distinguished by an additional input embedding for the frame index. Notably, even if only a single frame is generated, the task is still classified as multi-frame generation.
@@
    Token-Level Understanding: This task focuses on predicting properties of individual tokens. For example, in microscopic structures, tokens (atoms) possess atom-level properties such as partial charges, energies, and forces. To support token-level understanding, Uni-3DAR incorporates an additional prediction head on fine-grained structural tokens.
@@
    Structure-Level Understanding: This task captures global properties of 3D structures. In small molecules, for example, key properties include solubility, toxicity, and quantum mechanical attributes. To enable structure-level understanding, Uni-3DAR attaches a dedicated prediction head to the “[EoS]” (End of Structure) token. Given the scarcity of labeled data, prior works frequently employ unsupervised pretraining to enhance performance. Similarly, we can pretrain Uni-3DAR on large-scale unlabeled 3D structural data using the masked next-token prediction approach, enabling efficient fine-tuning for property prediction in alignment with existing models.
In Uni-3DAR, each token is assigned a distinct role, enabling the model to address four tasks within a unified framework. Specifically, generation (both single- and multi-frame) is performed on masked tokens, token-level understanding on fine-grained structural tokens, and structure-level understanding via the “[EoS]” token. Because these tasks operate independently on different tokens, they do not interfere with one another. As a result, Uni-3DAR unifies 3D structural generation and understanding in a single model and supports joint training across tasks.
Furthermore, the autoregressive nature of Uni-3DAR allows seamless integration of data from other modalities. This capability is particularly valuable because numerous 3D structural tasks inherently involve multi-modal inputs. For instance, protein folding commonly relies on protein sequences as conditioning inputs to generate corresponding 3D structures. Similarly, tasks such as predicting microscopic 3D structures from powder X-ray diffraction (PXRD) or nuclear magnetic resonance spectroscopy (NMR) signals benefit from a multi-modal, autoregressive approach. Sec. 3.2 further illustrates this approach through the example of PXRD-guided crystal structure prediction.
2.4 Implementation Details
We detail the implementation of microscopic 3D structure modeling, which is the primary focus of this study. Notably, apart from the fine-grained tokenization tailored for atoms, most other components can be seamlessly adapted to other 3D structures.
Tokenization Details
We always apply a random rotation before tokenization for data augmentation. For all experiments, we set the final tree level’s cell edge length to $c_{L-1}=0.24\text{\AA}$ and use a resolution of $c_{r}=0.01\text{\AA}$. The number of tree levels $L$ is chosen based on the system size. With these settings, the token count is approximately 10 times the number of atoms after applying both 2-level subtree compression and duplicated masked tokens. For example, on the QM9 dataset [25], we use $L=6$, resulting in roughly 160 tokens per structure (with an average of about 18 atoms), whereas using a full-size grid would yield $(2^{6})^{3}=262,144$ tokens. During training, we found that randomly dropping a small fraction of tokens from the final octree layer (using a ratio sampled from $[0,\gamma]$, where $\gamma$ is a hyperparameter) improves overall performance. We hypothesize that this noise injection enhances robustness in generating fine-grained tokens, thereby boosting final performance.
Model Architecture
We adopt a modern transformer decoder-only architecture [19]. Each layer consists of a uni-directional self-attention module and a SwiGLU [26] feed-forward network. We use a pre-norm design [27] with RMSNorm [28] for normalization. Unless otherwise specified, our default configuration is based on the GPT-2 model size, with 12 layers, an embedding dimension of 768, and a head dimension of 64, resulting in a model with approximately 90M parameters.
Input Embedding and Positional Encoding
For the $i$-th token, its content consists of a token type and in-cell positions, denoted as $(t_{i},\bm{e}_{i})\in\mathbb{N}^{4}$. We further define its tree level as $l_{i}\in\mathbb{N}$, its frame index as $f_{i}\in\mathbb{N}$ (if it is a multi-frame generation), and its cell position as $\bm{c}_{i}\in\mathbb{R}^{3}$. For octree and masked tokens, $\bm{c}_{i}$ is set to the center of the corresponding cell; for atom tokens, where each cell contains one atom, we use the atom’s precise coordinate for $\bm{c}_{i}$ to better represent its position. Then, $(t_{i},\bm{e}_{i},l_{i},f_{i})$ are converted into numerical features by summing their corresponding embedding layers, which then serve as the model’s input. Notably, we do not include any features from 2D graphs, such as bond information, ensuring that the proposed method can be applied to a wide range of 3D data. For pair-wise positional information, we simply apply RoPE-3D [29] to $\bm{c}_{i}$.
Generation Heads
The generation tasks are performed exclusively on the masked token. For octree tokens, since $\bm{e}_{i}$ is fixed to $N_{p}/2$, only $t_{i}$ needs to be predicted, which can be accomplished with a simple classification head. In contrast, for atom tokens, both $t_{i}$ and $\bm{e}_{i}$ must be predicted. We implement two solutions: autoregressive prediction and diffusion prediction. In both approaches, we first predict $t_{i}$. Then, for autoregressive prediction, we sequentially predict $(e_{i}^{0},e_{i}^{1},e_{i}^{2})$, using Alg. 1. For diffusion prediction, we adopt the token-level diffusion module from MAR [23] to predict $\bm{e}_{i}$ within each token. Since $\bm{e}_{i}$ originates from a continuous 3D space, employing diffusion to generate it is a reasonable choice. In our empirical experiments, both methods achieve similar performance (See Sec.3.8); therefore, by default, we use the more efficient autoregressive prediction. During inference, to balance quality and diversity, we first sample with a slightly higher temperature to enhance diversity, then select the top-$r$ samples ranked by cumulative autoregressive probabilities. Empirically, this approach outperforms low-temperature sampling in generating diverse yet high-quality samples.
Efficiency Optimizations
During training, we leverage FlashAttention [30] with bfloat16 to accelerate computation and reduce peak memory usage. Additionally, we employ sequence packing by concatenating tokens from multiple samples into a single large sequence, thereby eliminating the extra cost associated with traditional padding solutions. With these optimizations, Uni-3DAR can efficiently train on large systems with many atoms, such as proteins. During inference, we utilize a KV-cache to accelerate prediction. Furthermore, to accelerate masked next-token prediction, we generate tokens in pairs rather than one-by-one as in standard next-token prediction. Since the inputs for masked tokens are predetermined and do not require model prediction, we can pack two adjacent tokens together to increase GPU utilization.
1:  $y\leftarrow x$
2:  Initialize $preds\leftarrow\{\}$
3:  for $i\leftarrow 1$ to $n$ do
4:     $p\leftarrow pred\_heads[i](y)$
5:     Append $p$ to $preds$
7:  end for
8:  return  $preds$
@@ 3 Experiments
In this section, we perform real-world experiments to empirically validate the effectiveness of our proposed framework. Specifically, Table 1 summarizes key information across these tasks, illustrating that Uni-3DAR can address a diverse range of tasks with various data types within a unified framework.
Sec. 3.1 Molecule
Sec. 3.2 Crystal + PXRD
Sec. 3.3 Protein
Sec. 3.4 Protein + Molecule
Sec. 3.5 Molecule
Sec. 3.6 Polymer
@@
Dataset and Metric
Consistent with previous studies [11], we use the QM9 [31] and GEOM-DRUG [32] datasets for unconditional 3D molecular generation. QM9, a widely-used molecular machine learning benchmark, contains 130K small molecules with high-quality 3D conformations (up to 9 heavy atoms and 29 total atoms including hydrogens), split into training (100K), validation (18K), and test sets (13K). GEOM-DRUG, in contrast, features larger organic compounds containing up to 181 atoms (averaging 44.2 atoms across 5 types), covering approximately 37 million conformations for around 450K unique molecules. Following established protocols [11], we select the 30 lowest-energy conformations per molecule for training.
Model performance is evaluated based on chemical feasibility. Bond types (single, double, triple, or none) are inferred from molecular geometries using pairwise atomic distances and atom types. Metrics include Atom Stability (the fraction of atoms exhibiting correct valency), Molecule Stability (the percentage of molecules where all atoms are stable), validity (percentage of chemically valid molecules verified by RDKit), and uniqueness (percentage of unique compounds among generated molecules). Metrics are computed consistently using the evaluation code from previous studies [11].
Baselines and Implementation
We benchmark Uni-3DAR against established models, including G-SchNet [33], ENF [34], EDM [11] and its variants GDM [11], EDM-Bridge [35], GeoLDM [12], and UniGEM [36], which uses additional molecular properties to enhance generation performance.
Results
Table 2 clearly demonstrates Uni-3DAR’s superior performance, significantly surpassing all baseline models. On the QM9 dataset, Uni-3DAR achieves notable improvements in Molecule Stability and Validity, substantially outperforming the second-best method. These results highlight Uni-3DAR’s robust capability to generate high-quality molecules. Furthermore, even compared to UniGEM, which benefits from additional molecular property information, Uni-3DAR achieves superior performance using only 3D molecular data. This further underscores the efficacy and robustness of the proposed Uni-3DAR model.
Table 2: Performance comparison on unconditional 3D molecular generation. Results for UniGEM are marked with an asterisk (*) to indicate the use of additional molecular property information during training to enhance generation performance.
@@ Results
Atom Sta(%$\uparrow$ Mol Sta(%$\uparrow$ Valid(%$\uparrow$ V $\times$ U(%$\uparrow$ Atom Sta(%$\uparrow$ Valid(%$\uparrow$
@@
ENF [34] 85.0 4.9 40.2 39.4 - -
G-Schnet [33] 95.7 68.1 85.5 80.3 - -
GDM [11] 97.0 63.2 - - 75.0 90.8
GDM-AUG [11] 97.6 71.6 90.4 89.5 77.7 91.8
EDM [11] 98.7 82.0 91.9 90.7 81.3 92.6
EDM-Bridge [35] 98.8 84.6 92.0 90.7 82.4 92.8
GeoLDM [12] 98.9 89.4 93.8 92.7 84.4 99.3
UniGEM* [36] 99.0 89.8 95.0 93.2 85.1 98.4
@@
3.2 Crystal Generation
Tasks
  1. 1.
  2. 2.
  3. 3.
Dataset and Metric
We employ established datasets consistent with prior studies [37, 38, 39] for both training and evaluation purposes. Specifically, we employ the Carbon-24 dataset [40], containing 10,153 carbon-based structures with cells composed of 6 to 24 atoms. The MP-20 dataset [41], derived from the Materials Project [41], includes 45,231 stable inorganic materials representing a wide range of experimentally validated compounds, each containing up to 20 atoms per cell. Additionally, we use the more challenging MPTS-52 dataset, an extended version of MP-20, comprising 40,476 structures with up to 52 atoms per cell, organized by the earliest publication year. We follow the same data split strategy as outlined in previous work [38].
For assessing performance in CSP and PXRD-guided CSP tasks, we align our evaluation methodology with prior research [39]. We compute the top-1 match rate alongside the corresponding average root-mean-square error (RMSE) for matched structures. We employ StructureMatcher[42], using thresholds set to stol=0.5, angle_tol=10, and ltol=0.3, consistent with the methodology of previous studies [39].
Table 3: Results on de novo crystal generation. Baseline results are taken from Xie et al. [37]. For FlowMM, we report its result with 1000 integration steps.
@@ 3.2 Crystal Generation
Carbon-24 FTCP [43] 0.08 0.00 0.00 5.206 19.05
G-SchNet [44] 99.94 0.00 0.00 0.9427 1.320
P-G-SchNet [44] 48.39 0.00 0.00 1.533 134.7
CDVAE [37] 100.0 99.80 83.08 0.1407 0.2850
DiffCSP [38] 100.0 99.90 97.27 0.0805 0.0820
@@
MP-20 FTCP [43] 1.55 48.37 4.72 0.09 23.71 160.9 0.7363
G-SchNet [44] 99.65 75.96 38.33 99.57 3.034 42.09 0.6411
P-G-SchNet [44] 77.51 76.40 41.93 99.74 4.04 2.448 0.6234
CDVAE [37] 100.0 86.70 99.15 99.49 0.6875 0.2778 1.432
DiffCSP [38] 100.0 83.25 99.71 99.76 0.3502 0.1247 0.3398
FlowMM [39] 96.85 83.19 99.49 99.58 0.239 0.083
@@
CDVAE [37] 17.09 0.2969 5.34 0.2106 33.90 0.1045
DiffCSP [38] 17.54 0.2759 12.19 0.1786 51.49 0.0631
FlowMM [39] 23.47 0.4122 17.54 0.1726 61.39 0.0566
UniGenX [45] 27.09 0.2264 29.09 0.1256 63.88 0.0598
PXRDGEN [46] 68.68 0.0707

### v3

@@ 2.3 Unified 3D Generation and Understanding Framework
By integrating techniques in sections˜2.1 and 2.2, Uni-3DAR provides a unified framework for a wide range of 3D tasks (Figure˜3(b)). The model architecture assigns distinct roles to different token types, enabling it to handle four primary task categories individually or jointly:
@@
    Single-Frame Generation (sections˜3.1, 3.3, 3.4 and 3.6): Generating a single 3D structure, either unconditionally or conditioned on external modalities like text or chemical properties. This is accomplished using the masked tokens for autoregressive generation.
@@
    Multi-Frame Generation (sections˜3.2 and 3.5) Autoregressively producing a sequence of 3D structures (multiple octrees), such as a molecular dynamics trajectory, molecular docking, or pocket-based generation. Each frame is distinguished by a unique frame-index embedding.
@@
    Token-Level Understanding (section˜3.4) Predicting properties of local components (e.g., atomic forces or partial charges) by attaching a prediction head to the fine-grained tokens.
@@
    Structure-Level Understanding (section˜3.6): Predicting global properties of an entire structure (e.g., solubility, toxicity) via a prediction head on the final “[EoS]” token. This allows Uni-3DAR to be pre-trained on large-scale unlabeled 3D data and efficiently fine-tuned for downstream tasks.
This versatile design allows for seamless joint training across these diverse tasks. Each token type serves a clear purpose: masked tokens drive generation, fine-grained tokens facilitate local understanding, and the “[EoS]” token enables global understanding.
Furthermore, the autoregressive nature of Uni-3DAR inherently supports multi-modal conditioning, which is critical for many scientific applications. For instance, a protein’s amino acid sequence can guide the generation of its 3D fold. Similarly, experimental data like Powder X-ray Diffraction (PXRD) spectra can constrain the prediction of a crystal structure, a task we explore in section˜3.2.
@@ 3 Experiments
We conducted extensive experiments to validate Uni-3DAR across diverse benchmarks. This section summarizes the key findings; complete implementation details settings can be found in appendices˜B and C. To ensure a fair comparison against existing methods, we trained separate model for each benchmark. We defer the investigation of joint training to future work. Uni-3DAR is robust to hyper-parameters, requiring no significant tuning and using a consistent setting across all tasks.
@@
We assess Uni-3DAR on unconditional 3D molecular generation, a fundamental task challenged by the need to produce realistic conformations while accounting for molecular flexibility and diverse rotatable bonds. Our evaluation employs two standard benchmarks: QM9 (Ramakrishnan et al., 2014b), a dataset of small molecules with up to 29 atoms, and GEOM-DRUG (Axelrod and Gomez-Bombarelli, 2022), which contains larger, more complex drug-like compounds with up to 181 atoms. Following the established protocols of Hoogeboom et al. (2022), we report on key metrics including Atom Stability, Molecule Stability, chemical validity (as determined by RDKit), and uniqueness. Bond types are inferred from the generated geometries to evaluate chemical correctness.
As shown in table˜1, Uni-3DAR significantly outperforms all baseline models. On QM9, it achieves notable improvements in crucial metrics, reaching a Molecule Stability of $93.7\%$ and a Validity of $98.0\%$, substantially exceeding the second-best methods. These results underscore Uni-3DAR’s robust capability to generate high-quality, chemically valid molecules. Furthermore, Uni-3DAR surpasses UniGEM, a model that leverages additional molecular property information during training, using only 3D geometric data. This highlights the efficacy and robustness of our proposed model.
Table 1: Performance comparison on unconditional 3D molecular generation. Results for UniGEM are marked with an asterisk (*) to indicate the use of additional molecular property information during training.
@@
[…]
@@
ENF (Garcia Satorras et al., 2021) 85.0 4.9 40.2 39.4 - -
G-Schnet (Gebauer et al., 2022) 95.7 68.1 85.5 80.3 - -
GDM (Hoogeboom et al., 2022) 97.0 63.2 - - 75.0 90.8
GDM-AUG (Hoogeboom et al., 2022) 97.6 71.6 90.4 89.5 77.7 91.8
EDM (Hoogeboom et al., 2022) 98.7 82.0 91.9 90.7 81.3 92.6
EDM-Bridge (Wu et al., 2022) 98.8 84.6 92.0 90.7 82.4 92.8
GeoLDM (Xu et al., 2023b) 98.9 89.4 93.8 92.7 84.4 99.3
UniGEM* (Feng et al., 2024) 99.0 89.8 95.0 93.2 85.1 98.4
@@
Table 2: Results on de novo crystal generation. Baseline results are taken from Xie et al. (2021).
@@
Carbon-24 FTCP (Ren et al., 2021) 0.08 0.00 0.00 5.206 19.05
G-SchNet (Gebauer et al., 2019) 99.94 0.00 0.00 0.9427 1.320
P-G-SchNet (Gebauer et al., 2019) 48.39 0.00 0.00 1.533 134.7
CDVAE (Xie et al., 2021) 100.0 99.80 83.08 0.1407 0.2850
DiffCSP (Jiao et al., 2023) 100.0 99.90 97.27 0.0805 0.0820
@@
MP-20 FTCP (Ren et al., 2021) 1.55 48.37 4.72 0.09 23.71 160.9 0.7363
G-SchNet (Gebauer et al., 2019) 99.65 75.96 38.33 99.57 3.034 42.09 0.6411
P-G-SchNet (Gebauer et al., 2019) 77.51 76.40 41.93 99.74 4.04 2.448 0.6234
CDVAE (Xie et al., 2021) 100.0 86.70 99.15 99.49 0.6875 0.2778 1.432
DiffCSP (Jiao et al., 2023) 100.0 83.25 99.71 99.76 0.3502 0.1247 0.3398
FlowMM (Miller et al., 2024) 96.85 83.19 99.49 99.58 0.239 0.083
@@
CDVAE (Xie et al., 2021) 17.09 0.2969 5.34 0.2106 33.90 0.1045
DiffCSP (Jiao et al., 2023) 17.54 0.2759 12.19 0.1786 51.49 0.0631
FlowMM (Miller et al., 2024) 23.47 0.4122 17.54 0.1726 61.39 0.0566
UniGenX (Zhang et al., 2025) 27.09 0.2264 29.09 0.1256 63.88 0.0598
PXRDGEN (Li et al., 2024a) 68.68 0.0707